OpenStudy (mendicant_bias):

Doing a triple integral dealing with Divergence in Spherical coordinates, just want to make sure everything is right so far.

OpenStudy (mendicant_bias):

http://i.imgur.com/Ae4obGd.png

OpenStudy (mendicant_bias):

Just need to insert the Jacobian/integrating factor for Spherical coordinates in there, along with the 3x converted to spherical; but is the general setup right?

OpenStudy (anonymous):

You need to replace 3x by \[3 x=3 \cos (\theta ) \sin (\phi ) \]

OpenStudy (anonymous):

and \[ dxdydz= \rho ^2 \sin (\phi )d\phi d\theta d\rho \]

OpenStudy (mendicant_bias):

Alright, cool, but the setup is otherwise correct? I'm going to finish it now.

OpenStudy (mendicant_bias):

http://i.imgur.com/uZ1iD6o.png

OpenStudy (mendicant_bias):

http://i.imgur.com/kJTwPQR.png

OpenStudy (mendicant_bias):

I ended up getting 2pi, but the answer is 3pi.

OpenStudy (anonymous):

Sorry, there was a type \[ 3 x=3 \rho \cos (\theta ) \sin (\phi ) \]

OpenStudy (anonymous):

\[ \int_0^{\frac{\pi }{2}} \left(\int_0^{\frac{\pi }{2}} \left(\int_0^2 3 \rho ^3 \cos (\theta ) \sin ^2(\phi ) \, d\rho \right) \, d\theta \right) \, d\phi =3 \pi \]